OpenStudy (anonymous):

Line m has no y-intercept, and its x-intercept is (3, 0). Line n has no x-intercept, and its y-intercept is (0, -4). The equation of line m is , and the equation of line n is

OpenStudy (decentnabeel):

the equation is.. y-y1=m(x-x1) first find the m m is the slope..

OpenStudy (decentnabeel):

do you know slope @Kaelyn78

OpenStudy (anonymous):

no they don't give the slope

OpenStudy (decentnabeel):

slope=(y2-y1)/(x2-x1) (3,0)(0,-4) x1=3 y1=0 x2=0 y2=-4 so find it slope

OpenStudy (anonymous):

line m is vertical and line n is horizontal

OpenStudy (anonymous):

Vertical lines have the equation: x = x-intercept Horizontal lines have the equation: y = y-intercept

OpenStudy (decentnabeel):

m=-4/3 so we use that (3,0) the equation is.. y-0=-4/3(x-3) y=-4/3x+4

OpenStudy (anonymous):

@DecentNabeel there will be two different equations, one for each line. I think you did the equation for the line that passes through the given points.

OpenStudy (anonymous):

I'm looking for two different equations, one for line m, another for line n

OpenStudy (anonymous):

For m, since it's vertical the equation is x = x-intercept. The x-intercept is 3, so the equation is x = 3.

OpenStudy (anonymous):

follow the same pattern for n, but since it's a horizontal line it's y = y-intercept

OpenStudy (anonymous):

so y=-3?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

yw